# Role of long-range dipolar interactions in the simulation of the properties of polar crystals using effective atomic potentials

### v1

@@ I Introduction
For decades, empirical and semi-empirical models have been fundamental in the atomistic modeling of materials, providing an alternative framework even prior to first-principles methods. Early approaches relied on interatomic potentials such as Lennard-Jones [1], Stillinger-Weber [2], and Tersoff [3] potentials, which successfully described properties of simple materials. Rapidly they improved and became more sophisticated [4] and today a wide range of strategies, depending on the specific material system, can be found in the literature [5, 6] providing highly accurate predictions of material properties. With advances in computing power and electronic‑structure theory [7, 8, 9, 10], the community increasingly turned to first‑principles simulations for greater predictive accuracy. However, aided by modern parameter fitting algorithms, high-quality reference data, and new perspectives, the atomistic modeling of materials is now enjoying a resurgence, allowing simulations of longer length and time scales as well as including operating conditions, such as the application of external electric fields or mechanical constraints [11]. Atomistic models directly fitted on first-principles data (and sometimes referred to as “second-principles”) now offer a versatile framework that allows researchers to create custom models for a wide range of materials ranging from metals and semiconductors to insulators [12, 13, 14, 15]. Modern approaches include pioneering lattice effective Hamiltonians [16, 17], related effective atomic potentials [18, 19, 20, 21], phase-fields methodologies [22], shell- and bond-valence-models [23, 24] or recent machine-learning interatomic potentials, including emerging universal potentials [25].
@@
Recently, with the advent of machine learning, atomic potentials have indeed experienced a revolution and are rapidly becoming a standard complementary tool in computational materials science [12, 13, 26, 27, 14, 15] to extend the predictive power of first-principles simulations in a multiscale context. Approaches such as Gaussian process regression [28], kernel methods [27] and neural networks [26, 29] have democratized the generation of accurate interatomic models for many different materials. These models are typically trained on extensive datasets generated from first-principles calculations without the inclusion of any extra terms and relying on the configuration itself rather than distortions with respect to a high-symmetry reference structure. As a result, long-range interactions are often neglected. Although these models are typically assumed to depend on the local atomic environment at first sight, the actual range of the interactions is dependent and inherently determined by the size of the training set used to develop the model. Therefore, the correct description of the nonanalytic behavior of the interatomic force constants in the long-wavelength limit presents a challenge as it originates from the power-law behavior of the Coulombic interaction and cannot be easily captured by spatially truncated training sets. Moreover, the absence of a reference structure inherently complicates the analytical inclusion of electrostatic interactions, which are often neglected in AI models beyond the spatial range covered by their training set. Novel approaches are trying to circumvent this issue including charge equilibration schemes [30], the on-the-fly prediction of maximally localized Wannier function centers [31, 32] or a recently proposed modification of the model without the need for additional first-principles calculations or retraining discussed in Ref. [33]. However, these tools have not yet become common practice, and most of the machine-learned interatomic potentials rely on the nearsightedness of interatomic interactions neglecting such long range part of the interactions beyond the range defined by the training set which in most cases turns out to be computed on a 22 supercell.
@@
In this work, we aim to clarify the concrete impact of neglecting the long-range part of the dipole-dipole interactions in computational models and to identify the scenarios where incorporating corrections for an accurate description becomes crucial. By focusing on the prototypical ferroelectric material BaTiO3 in which dipolar interactions are expected to be key [34, 35], we systematically explore the influence of the dipole-dipole interactions on the phonon dispersion curves and atomic distortion patterns. On the one hand, we demonstrate that the inclusion of the long-range part of dipolar interactions is not always essential for an overall accurate description of the system, including the ferroelectric instability and its characteristic chain-like character, which supports the good performance of AI-based models even in polar materials [36]. On the other hand, we also point out specific cases where neglecting the long-range part of the dipole-dipole interaction can lead to significant artifacts in the phonon spectrum, resulting in inaccurate descriptions of the energy landscape associated with certain atomic displacements and the emergence of spurious minima.
@@ II Methodology
To assess the impact of the long-range character of dipolar interactions on the lattice dynamics of BaTiO3, we rely here on a second-principles effective atomic potential (EAP) method, as implemented in the multibinit software. The method relies on a Taylor expansion of the the Born-Oppenheimer potential energy surface (PES) in terms of structural degrees of freedom around the cubic perovskite structure taken as reference. The expansion is made in terms of individual atomic displacements and homogeneous strain degrees of freedom and includes phonon, strain and strain-phonon energy terms at both harmonic and anharmonic levels. The coefficients of harmonic terms corresponds to second-energy derivatives which are straightforwardly determined from density-functional perturbation theory, so that the model keeps first-principles accuracy at the harmonic level. Then, most relevant anharmonic terms, which capture key deviations from harmonic behavior, are selected and their coefficients fitted in order to reproduce the energies, forces, and stresses as obtained from DFT calculations for a representative set of atomic configurations. The reference model used here is a slight revision of the one reported in Ref. [37], only including refitted anharmonic terms. Parameters and a validity assessment passport are provided in Appendix.
@@
Remarkably, the model derived with a 2×2×2 -point mesh, which excludes the analytical dipole-dipole correction, is of special interest, as it mimics the typical interaction range used in the training sets of AI models which in turn defines the maximum range of interactions these models typically incorporate. At the other extreme, the model with 8×8×8 -point mesh, including the analytical dipole-dipole correction can be considered as the exact model to be taken as reference.
@@ III Results
This establishes for instance, that if a BaTiO3 model is built using AI methods with a 222 supercell for the training set, and so limiting implicitly the interaction range to 2 unit cells (equivalently to our 222 model without dipole-dipole), the overall description of the material will remain largely accurate (the ferroelectric instability and its chain-like character [39] are preserved), except for the specific LO phonon branches affected by the missing long-range part of dipolar interactions, particularly near . This discrepancy is unlikely to significantly impact the material’s dynamics and phase transitions, as the phonon density of states remains well reproduced and the number of poorly described phonon bands is marginal. However, as it will be emphasized later, it may introduce artificial minima in the energy landscape, potentially affecting specific analyses. It must also be noted that, although the present discussion might have some general character, the specific 222 range is a priori material dependent and cannot be taken for granted in a general case.
It is important at this stage to clarify the distinction between the plotted phonon dispersion curves and the information within the model. Programs like Phonopy [44] can include the contribution of dipole-dipole interactions on top of the information provided by the M2 model, thereby mimicking the phonon dispersion curves of M2+Dip model plotted in Fig.3(f).
This is indeed a common practice in the literature [45, 46, 36]. However, this correction is artificially made at the level of the post-processing while it is not present in the independent use of the model creating the illusion that the model phonons correspond to those plotted on Fig.3(f), while in reality, the model remains at the level of Fig. 3(i).
@@
To point out the main and concrete deficiencies arising from the exclusion of the long-range component of the dipole-dipole interactions, we can condense the LO phonon at u.c.-1 from the spurious unstable branch predicted by the M2 model (Fig. 4a) As shown in Fig. 5, the condensation of such LO modes induce head to head and tail to tail domains that remain metastable due to the wrong description of the phonon branch. Clearly, this polarization state is unphysical as the divergence of the polarization generates bound charges on the material that enormously increase the electrostatic energy of the system. Similarly, when condensing together two of these LO phonon modes directed along and , we can stabilize a lattice of center-convergent and center-divergent vortices demonstrating that more complex polar phases can also suffer from this spurious behavior. Properly including the long-range part of the dipole-dipole interactions, as discussed earlier, stabilizes this phonon branch and corrects the issue, leading to the suppression of the associated fake metastable states. As shown in Fig. 5(c), the energy landscape as a function of the phonon mode amplitude changes dramatically with the inclusion of the long-range part of the dipole-dipole interactions. The original double-well potential is replaced by a steeply rising energy surface, preventing the formation of bound charges and restoring a proper description of the material. In conclusion, it is important to note that as soon as the longitudinal optical phonon modes are involved in the atomic displacements of the material under study, the proper treatment of the long-range part of the dipole-dipole interactions remains crucial to avoid producing fake minima. This might become relevant when studying topological textures, dynamics of large supercells or thin films and superlattices where depolarizing fields play an important role.
@@ IV Conclusions
In summary, our results demonstrate that the overall phonon dispersion curves are well captured even when the interactions are limited to a relatively small range which explains the success of AI-based models in describing material properties, where the range of interactions is inherently limited by the supercell size used to generate the training set. Particularly surprising is the fact that the ferroelectric instability is very well described even in the case of BaTiO3, where long-range electrostatic interactions were thought to be crucial for its ferroelectric character. Our results show, however, that while these dipolar interactions are crucial, it is the short-range part of them that appears to play a key role in driving this effect.

### v2

@@ I Introduction
For decades, empirical and semi-empirical models have been fundamental in the atomistic modeling of materials, providing an alternative framework even prior to first-principles methods. Early approaches relied on interatomic potentials such as Lennard-Jones [1], Stillinger-Weber [2], and Tersoff [3] potentials, which successfully described properties of simple materials. Rapidly, they improved and became more sophisticated [4] and today a wide range of strategies, depending on the specific material system, can be found in the literature [5, 6] providing highly accurate predictions of material properties. With advances in computing power and electronic‑structure theory [7, 8, 9, 10], the community increasingly turned to first‑principles simulations for greater predictive accuracy. However, aided by modern parameter fitting algorithms, high-quality reference data, and new perspectives, the atomistic modeling of materials is now enjoying a resurgence, allowing simulations of longer length and time scales as well as including operating conditions, such as the application of external electric fields or mechanical constraints [11]. Atomistic models directly fitted on first-principles data (and sometimes referred to as “second-principles”) now offer a versatile framework that allows researchers to create custom models for a wide range of materials ranging from metals and semiconductors to insulators [12, 13, 14, 15]. Modern approaches include pioneering lattice effective Hamiltonians [16, 17], related effective atomic potentials [18, 19, 20, 21], phase-fields methodologies [22], shell- and bond-valence-models [23, 24], reactive force fields [25] and their application to ferroelectrics [26, 27, 28] or recent machine-learning interatomic potentials, including emerging universal potentials [29].
@@
Recently, with the advent of machine learning, atomic potentials have indeed experienced a revolution, and are rapidly becoming a standard complementary tool in computational materials science [12, 13, 30, 31, 14, 15] to extend the predictive power of first-principles simulations in a multiscale context. Approaches such as Gaussian process regression [32], kernel methods [31] and neural networks [30, 33] have democratized the generation of accurate interatomic models for many different materials. These models are typically trained on extensive datasets generated from first-principles calculations without the inclusion of any extra terms and relying on the configuration itself rather than distortions with respect to a high-symmetry reference structure. As a result, long-range interactions are often neglected. Although these models are typically assumed to depend on the local atomic environment at first sight, the actual range of the interactions is dependent and inherently determined by the size of the training set used to develop the model. Therefore, the correct description of the nonanalytic behavior of the interatomic force constants in the long-wavelength limit presents a challenge as it originates from the power-law behavior of the Coulombic interaction and cannot be easily captured by spatially truncated training sets. Moreover, the absence of a reference structure inherently complicates the analytical inclusion of electrostatic interactions, which are often neglected in AI models beyond the spatial range covered by their training set. Novel approaches are trying to circumvent this issue including charge equilibration schemes [34], the on-the-fly prediction of maximally localized Wannier function centers [35, 36] or a recently proposed modification of the model without the need for additional first-principles calculations or retraining discussed in Ref. [37]. However, these tools have not yet become common practice, and most of the machine-learned interatomic potentials rely on the nearsightedness of interatomic interactions neglecting such long range part of the interactions beyond the range defined by the training set which in most cases turns out to be computed on a 22 supercell.
@@
In this work, we aim to clarify the concrete impact of neglecting the long-range part of the dipole-dipole interactions in computational models and to identify the scenarios where incorporating corrections for an accurate description becomes crucial. By focusing on the prototypical ferroelectric material BaTiO3 in which dipolar interactions are expected to be key [38, 39], we systematically explore the influence of the dipole-dipole interactions on the phonon dispersion curves and atomic distortion patterns. On the one hand, we demonstrate that the inclusion of the long-range part of dipolar interactions is not always essential for an overall accurate description of the system, including the ferroelectric instability and its characteristic chain-like character, which supports the good performance of AI-based models even in polar materials [40]. On the other hand, we also point out specific cases where neglecting the long-range part of the dipole-dipole interaction can lead to significant artifacts in the phonon spectrum, resulting in inaccurate descriptions of the energy landscape associated with certain atomic displacements and the emergence of spurious phases.
@@ II Methodology
To assess the impact of the long-range character of dipolar interactions on the lattice dynamics of BaTiO3, we rely here on a second-principles effective atomic potential (EAP) method, as implemented in the multibinit software. The method relies on a Taylor expansion of the Born-Oppenheimer potential energy surface (PES) in terms of structural degrees of freedom around the cubic perovskite structure taken as reference. Although lower symmetry phases like the , , or could also be considered as proper references for constructing the model, the choice of the cubic parent phase follows previous similar studies in perovskites based on Taylor expansion [16, 18, 11] and is the most sensible choice to access the whole phase diagram. The Taylor expansion around the reference geometry is made in terms of individual atomic displacements and homogeneous strain degrees of freedom and includes phonon, strain and strain-phonon energy terms at both harmonic and anharmonic levels. The coefficients of harmonic terms corresponds to second-energy derivatives which are straightforwardly determined from density-functional perturbation theory, so that the model keeps first-principles accuracy at the harmonic level. Then, most relevant anharmonic terms, which capture key deviations from harmonic behavior, are selected and their coefficients fitted in order to reproduce the energies, forces, and stresses as obtained from DFT calculations for a representative set of atomic configurations. The reference model used here is a slight revision of the one reported in Ref. [41], only including refitted anharmonic terms. Parameters and a validity assessment passport are provided in Appendix.
@@
Remarkably, the model derived with a 2×2×2 -point mesh, which excludes the analytical dipole-dipole correction, is of special interest, as it mimics the typical interaction range used in the training sets of AI models, which in turn defines the maximum range of interactions these models incorporate. At the other extreme, the model with 8×8×8 -point mesh, including the analytical dipole-dipole correction can be considered as the exact model to be taken as reference.
@@ III Results
This establishes for instance, that if a BaTiO3 model is built using AI methods with a 222 supercell for the training set, and so limiting implicitly the interaction range to 2 unit cells (equivalently to our 222 model without dipole-dipole), the overall description of the material will remain largely accurate (the ferroelectric instability and its chain-like character [46] are preserved), except for the specific LO phonon branches affected by the missing long-range part of dipolar interactions, particularly near . This discrepancy is unlikely to significantly impact the material’s dynamics and phase transitions, as the phonon density of states remains well reproduced and the number of poorly described phonon bands is marginal. This is indeed confirmed in Fig. 4 where it appears that the phase transition sequence and hysteresis loops reproduced by M2 and M8+Dip models are very closely similar which explains why AI-based models limited to short-range interactions have been shown to successfully predict phase diagrams and ferroelectric transitions in complex materials [40, 51, 52, 53, 54]. However, as it will be emphasized later, the improper treatment of the long-range part of dipolar interactions may introduce artificial discrepancies in the energy landscape, potentially affecting specific analyses.
It must also be noted that, although the present discussion might have some general character, the specific 222 range is a priori material dependent and cannot be taken for granted in a general case.
It is important at this stage to clarify the distinction between the plotted phonon dispersion curves and the information within the model. Programs like Phonopy [55] can include the contribution of dipole-dipole interactions on top of the information provided by the M2 model, thereby mimicking the phonon dispersion curves of M2+Dip model plotted in Fig.3(f). This is indeed a common practice in the literature [53, 56, 40]. However, this correction is artificially made at the level of the post-processing while it is not present in the independent use of the model creating the illusion that the model phonons correspond to those plotted on Fig.3(f), while in reality, the model remains at the level of Fig. 3(i).
@@
To point out the main and concrete deficiencies arising from the exclusion of the long-range component of the dipole-dipole interactions, we can condense the LO phonon at u.c.-1 from the spurious unstable branch predicted by the M2 model (Fig. 5a) As shown in Fig. 6, the condensation of such LO modes induce head-to-head and tail-to-tail domains that remain metastable due to the wrong description of the phonon branch. Clearly, this polarization state is unphysical as the divergence of the polarization generates bound charges on the material that enormously increase the electrostatic energy of the system. Similarly, when condensing together two of these LO phonon modes directed along and , we can stabilize a lattice of center-convergent and center-divergent vortices demonstrating that more complex polar phases can also suffer from this spurious behavior. As shown in Fig. 6, the energy lowering associated to such fake metastable phase is not negligible ( meV/f.u.) and comparable to that of the phase ( meV/f.u.).
Properly including the long-range part of the dipole-dipole interactions, as discussed earlier, stabilizes the problematic phonon LO branch and corrects the issue, leading to the suppression of the related fake low-energy states. As shown in Fig. 6(c), the energy landscape as a function of the phonon mode amplitude changes dramatically with the inclusion of the long-range part of the dipole-dipole interactions. The original double-well potential is replaced by a steeply rising energy surface, preventing the formation of bound charges and restoring a proper description of the material.
In the present examples, the apparent minima observed in Fig. 6(a) and (c) when limiting the range of interactions, are in fact saddle points of the Born-Oppenheimer energy surface once full structural relaxation without symmetry constraints is allowed. This is confirmed by finite-temperature simulations using the M2 models, where the associated atomic configurations are unstable and do not persist. This result is consistent with the absence of report of such metastable phases in previous AI models restricted to short-range interactions. But we cannot exclude the presence of eventual spurious local minima in other cases and, at least, our Fig. 6 demonstrates that models limited to short-range interactions will significantly over-stabilize head-to-head or tail-to-tail polar configurations, which will unavoidably affect phase competition.
In short, it is important to know that as soon as the longitudinal optical phonon modes are involved in the atomic displacements of the material under study, the proper treatment of the long-range part of the dipole-dipole interactions remains crucial to properly quantify the energetics. This might become relevant when studying topological textures, dynamics of large supercells or thin films and superlattices where depolarizing fields play an important role.
@@ IV Conclusions
In summary, our results demonstrate that the overall phonon dispersion curves as well as structural phase transitions with temperature and hysteresis loops are well captured even when the interactions are limited to a relatively small range which explains the success of AI-based models in describing material properties, where the range of interactions is inherently limited by the supercell size used to generate the training set. Particularly surprising is the fact that the ferroelectric instability is very well described even in the case of BaTiO3, where long-range electrostatic interactions were thought to be crucial for its ferroelectric character. Our results show, however, that while these dipolar interactions are crucial, it is the short-range part of them that appears to play a key role in driving this effect.
@@ Acknowledgements.
- van Duin et al. [2001] A. C. T. van Duin, S. Dasgupta, F. Lorant, and W. A. Goddard, ReaxFF: a reactive force field for hydrocarbons, The Journal of Physical Chemistry A 105, 9396 (2001).
- Akbarian et al. [2019] D. Akbarian, D. E. Yilmaz, Y. Cao, P. Ganesh, I. Dabo, J. Munro, R. Van Ginhoven, and A. C. T. van Duin, Understanding the influence of defects and surface chemistry on ferroelectric switching: a reaxFF investigation of BaTiO3, Phys. Chem. Chem. Phys. 21, 18240 (2019).
- Kelley et al. [2022] K. P. Kelley, A. N. Morozovska, E. A. Eliseev, V. Sharma, D. E. Yilmaz, A. C. T. van Duin, P. Ganesh, A. Borisevich, S. Jesse, P. Maksymovych, N. Balke, S. V. Kalinin, and R. K. Vasudevan, Oxygen vacancy injection as a pathway to enhancing electromechanical response in ferroelectrics, Advanced Materials 34, 2106426 (2022).
- Dhakane et al. [2023] A. Dhakane, T. Xie, D. E. Yilmaz, A. C. van Duin, B. G. Sumpter, and P. Ganesh, A graph dynamical neural network approach for decoding dynamical states in ferroelectrics., Carbon Trends 11, 100264 (2023).
@@
- Fletcher [2000] R. Fletcher, Practical methods of optimization (John Wiley & Sons, 2000).
- Duane et al. [1987] S. Duane, A. D. Kennedy, B. J. Pendleton, and D. Roweth, Hybrid monte carlo, Physics letters B 195, 216 (1987).
- Betancourt [2017] M. Betancourt, A conceptual introduction to Hamiltonian Monte Carlo, arXiv preprint arXiv:1701.02434 10.48550/arXiv.1701.02434 (2017).
@@
- Gigli et al. [2024] L. Gigli, A. Goscinski, M. Ceriotti, and G. A. Tribello, Modeling the ferroelectric phase transition in barium titanate with dft accuracy and converged sampling, Phys. Rev. B 110, 024101 (2024).
- Zhang and Chen [2025] C. Zhang and X. Chen, FerroAI: a deep learning model for predicting phase diagrams of ferroelectric materials, npj Computational Materials 11, 282 (2025).
- Feng et al. [2025] N. Feng, H.-C. Thong, K. Wang, K. Bi, and B. Xu, Atomic-scale polarization switching driven by terahertz light in ferroelectric , Phys. Rev. B 111, 024305 (2025).